OpenStudy (anonymous):

What is the value of g[f(3)] for the functions f(x) = 3x – 5 and g(x) = x + 1? (Type only the value of your final answer, no variables.) f(x) = 3(3) – 5 and g(x) = x + 1? =5

OpenStudy (bibby):

I got the same answer. First evaluate f(3) = 3*3 - 5 = 9 - 5 = 4 then g(4) = 4 + 1 = 5

OpenStudy (anonymous):

yep that's what I did, thank you!